Suppose a triangle is plotted on a coordinate plane with vertices located at ( -9, 3),(-3, -9) and (13, -1). What is the perimeter of the triangle?
Answer #1

of the given taiangle The vertices AC-9,3) B(-3,-9), ((13,-1) AB = 22-27 (429) V €3-6-91) 7 & 9-3) = √(3 + 9) + E2 / 64627 =